OpenStudy (anonymous):

2^8 x 4^6 x 8^8 please give the answers as an index number

OpenStudy (anonymous):

(2)^8(2^2)^6(2^3)^8 (2^8)(2^12)(2^24) 2^44

OpenStudy (anonymous):

is that answer equal to 4^22?

OpenStudy (anonymous):

Yes...since 2^44= 2^(2x22) = 4^22